# Lifetime measurement of the muonic atom of enriched Si isotopes

### v1

@@ VI Conclusion
The present experiment establishes a method for measuring the lifetimes of the muonic atom at MLF in J-PARC, opening up a new possibility for systematic investigations of capture rates. The re-evaluation of the muon capture rates with the Primakoff and Goulard-Primakoff formulas will provide better estimation in general-purpose simulation codes such as GEANT4 [25], PHITS [27], and FULUKA [28]. The systematic study of the capture rate will contribute to a further understanding of the quenching of and of the nuclear medium.
@@ Acknowledgements.
We are grateful for the valuable discussions with Dr. T. Y. Saito (RIKEN), Dr. H. Iwamoto (JAEA), and Dr. T. Miyagi (Univ. Tsukuba). The muon experiment at the Materials and Life Science Experimental Facility of the J-PARC was performed under a user program (Proposal No. 2024MI01). The present work is partially supported by the JSPS Grant-in-Aid under Grant Nos. JP19H05664, JP22K20372, JP23H04526, JP23H01845, JP23K01845, JP23K03426, JP24H00073, JP24K00647, and JP24K17057. R.M. is supported by the Force-front Physics and Mathematics Program in Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) program, and the JSR Fellowship from the University of Tokyo. T.N. acknowledges the RIKEN Special Postdoctoral Researcher Program. The numerical calculations were pertially performed on cluster computers at the RIKEN iTHEMS program.References
@@
- Iwamoto et al. [2024] H. Iwamoto, M. Niikura, and R. Mizuno, submitted to Phys. Rev. C (2024).
@@
- Mizuno et al. [2024] R. Mizuno, M. Niikura, T. Y. Saito, T. Matsuzaki, and others, to be submitted (2024).

### v2

@@ VI Conclusion
The present experiment establishes a method for measuring the lifetimes of the muonic atom at MLF in J-PARC, opening up a new possibility for systematic investigations of capture rates. The re-evaluation of the muon capture rates with the Primakoff and Goulard-Primakoff formulas will provide better estimation in general-purpose simulation codes such as GEANT4 [27], PHITS [29], and FLUKA [30]. The systematic study of the capture rate will contribute to a further understanding of the quenching of and of the nuclear medium.
@@ Acknowledgements.
We are grateful for the valuable discussions with Dr. T. Y. Saito (RIKEN), Dr. H. Iwamoto (JAEA), and Dr. T. Miyagi (Univ. Tsukuba). The muon experiment at the Materials and Life Science Experimental Facility of the J-PARC was performed under a user program (Proposal No. 2024MI01). The present work is partially supported by the JSPS Grant-in-Aid under Grant Nos. JP19H05664, JP22K20372, JP23H04526, JP23H01845, JP23K01845, JP23K03426, JP24H00073, JP24K00647, and JP24K17057. R.M. is supported by the Force-front Physics and Mathematics Program in Drive Transformation (FoPM), a World-leading Innovative Graduate Study (WINGS) program, and the JSR Fellowship from the University of Tokyo. T.N. acknowledges the RIKEN Special Postdoctoral Researcher Program. The numerical calculations were partially performed on cluster computers at the RIKEN iTHEMS program.References
@@
- Iwamoto et al. [2025] H. Iwamoto, M. Niikura, and R. Mizuno, Phys. Rev. C. 111, 034614 (2025).
@@
- Mizuno et al. [2025] R. Mizuno, M. Niikura, T. Y. Saito, T. Matsuzaki, S. Abe, H. Fukuda, M. Hashimoto, A. Hillier, K. Ishida, N. Kawamura, S. Kawase, T. Kawata, K. Kitafuji, F. Minato, M. Oishi, A. Sato, K. Shimomura, P. Strasser, S. Takeshita, D. Tomono, and Y. Watanabe, arXiv:2507.19753 [nucl-ex] (2025).
@@
- Wakata et al. [2025] M. Wakata, S. Akamatsu, T. Fujiie, T. Furuyama, L. Hara, Y. Ishikawa, T. Ito, T. Kikuchi, T. Mibe, S. Ozaki, M. Yokomizo, and J. Murata, Prog. Theor. Exp. Phys. 2025, 023H01 (2025).
- Takeshita et al. [2024] S. Takeshita, I. Umegaki, M. Tampo, P. Strasser, Y. Ikedo, A. Koda, W. Higemoto, T. Yuasa, N. Kawamura, Y. Miyake, and K. Shimomura, Interactions 245, 1 (2024).